In a town, 53% of the residents voted for Candidate F. If we randomly survey 4 residents, what is the probability that at least one of them did not vote for Candidate F?
Answer by ewatrrr(24785)


Hi
 .53 voted for candidate F
Chose 4 peopele:
P(at least one did NOT vote for F) = 1 - P(all voted for F) = 1-.47^4
 P = 1- .47^4 = .9512
Wish You the Best in your Studies.